A pilot plant has produced metallurgical batches which are summarized as follows: Low strength High... 1 answer below »


A pilot plant has produced metallurgical batches which are summarized as follows:


Low strength


High strength


Low in impurities


2


27


High in impurities


12


4


If these results are representative of full-scale production, find estimated probabilities that a production batch will be:


i) low in impurities


ii) high strength


iii) both high in impurities and high strength


iv) both high in impurities and low strength​


1
Expert's answer
2021-05-28T05:15:55-0400

Total number of results = 2+ 27+12+4 =45


(i) P(low in impurities ) = "\dfrac{2+27}{45}=\dfrac{29}{45}=0.645"


(ii) P(high strength ) = "\dfrac{27+4}{45}=\dfrac{31}{45}=0.689"


(iii) P(both high in impurities and high strength ) = "\dfrac{4}{45}=0.0889"


(iv) P(both high in impurities and low strength) = "\dfrac{12}{45}=0.2667"